# When Fusion Helps and When It Breaks: View-Aligned Robustness in Same-Source Financial Imaging

### v1

@@ 6.2 Adversarial Robustness
Protocol and summary metrics.
@@
Summary of findings.
We find (i) rapid degradation under tiny budgets, (ii) strong view asymmetry, and (iii) a clear robustness advantage of late fusion under view-constrained attacks, while robustness under joint (two-view) attacks remains challenging.
@@
Overall vulnerability under tiny $\ell_{\infty}$ budgets.
@@
View sensitivity and view-aligned evaluation.
Using the view–channel mapping in Sec. 3.3, we compare (i) single-view models attacked on their sole channel and (ii) two-view models attacked on the corresponding view channel (Sec. 4.2). The results reveal strong view dependence. In particular, ResNet18-P exhibits extreme sensitivity when the indicator view is perturbed: very small budgets can already drive MCC close to the lower bound, suggesting a near-consistent label-flip behavior rather than a gradual attenuation of predictive signal. In contrast, attacking the OHLCV view typically leads to a more gradual deterioration, highlighting asymmetric fragility across views.
Late fusion improves robustness under view-constrained attacks.
A consistent trend in the two-view setting is that dual-branch late fusion (*-cons) degrades more gracefully than early fusion (channel stacking) under view-constrained attacks. When only one view is perturbed, the unperturbed branch can still provide usable evidence, which mitigates worst-case collapse and yields a flatter robustness curve over $E$. This robustness advantage is especially visible when attacking the indicator channel, where late fusion maintains substantially higher MCC than early fusion over most of the budget range.
Joint perturbations remain challenging.
@@
Consistency regularization: secondary but non-negligible effects.
Within late-fusion models, increasing the consistency weight $\lambda$ can further reduce degradation under view-constrained attacks in some settings, but the dominant robustness gain comes from the architectural shift from early fusion to late fusion.
Model selection note: while $\lambda=0.5$ is the clean-optimal setting at $\tau=0.006$ (Sec. 6.1), we report late-fusion robustness summaries with $\lambda=1$ by default because it yields slightly better robustness (including the $2/255$ stress test) at a small cost in clean MCC. We therefore interpret consistency as a refinement on top of the late-fusion design rather than the primary driver.
@@
Table 3: PGD robustness summary at $\tau=0.006$ (min_move). We report MCC at two representative budgets ($0.25/255$ and $1/255$) and the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets $\{0.25,0.5,0.75,1\}/255$ (Eq. (9)). As an additional stress test, results at $\epsilon_{\mathrm{adv}}=2/255$ are reported in Appendix E.
@@
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@1/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$
Lite-CNN ohlcv on ohlcv 0.109 0.054 -0.088 -0.197
Lite-CNN indic on indic 0.079 0.024 -0.186 -0.265
Lite-CNN both (early) on ohlcv (ch0) 0.089 0.030 -0.104 -0.193
Lite-CNN both (early) on indic (ch1) 0.089 0.033 -0.094 -0.183
Lite-CNN both (early) on both 0.089 -0.014 -0.317 -0.406
Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.121 0.081 -0.029 -0.150
Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.121 0.113 0.084 -0.037
Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.121 0.071 -0.072 -0.192
ResNet18-P ohlcv on ohlcv 0.019 -0.443 -0.952 -0.971
ResNet18-P indic on indic 0.082 -0.999 -1.000 -1.082
ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.095 -0.211 -0.774 -0.870
ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.095 -0.959 -1.000 -1.095
ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.095 -0.983 -1.000 -1.095
Stress test at $\epsilon_{\mathrm{adv}}=2/255$ (not in the main grid).
Although our main evaluation grid stops at $1/255$, we additionally evaluate a larger budget as a stress test. Notably, under the indic-only attack (ch1), Lite-CNN-Cons maintains positive mean MCC at $2/255$ for both $\lambda=0.5$ and $\lambda=1$, with $\lambda=1$ being slightly stronger (Appendix E).
Branch diagnostics for late-fusion two-view models (clean test, $\tau=0.006$).
To interpret robustness trends for *-cons models, we inspect clean-test branch diagnostics that summarize cross-branch alignment (agree_ab, sym_kl with T=2T{=}2) and branch-wise versus fused MCC (Table 4; full results in Appendix F). For Lite-CNN-Cons, enabling consistency ($\lambda>0$) sharply increases alignment (higher agree_ab, much lower sym_kl), yet branch-wise MCC remains near zero while fused-head MCC stays substantially higher. This indicates that predictive signal is expressed primarily through the feature-level late-fusion pathway rather than strong standalone branch classifiers. For ResNet18P-Cons, larger $\lambda$ likewise improves alignment and is associated with higher fused MCC, but variance remains notable and robustness failures under indicator-view attacks persist.
[…]
Model $\boldsymbol{\lambda}$ agree_ab sym_kl mcc_a / mcc_b mcc_fuse
Lite-CNN-Cons 0 $0.275090\pm 0.244057$ $164898\pm 0.125751 0.0$$019377\pm 0.061519\ /\ 0.005906\pm 0.051856$ $0.127459\pm 0.035946$
Lite-CNN-Cons 0.5 $830197\pm 0.185333 0.0$$00195\pm 0.000098 0.0$$001488\pm 0.053262\ /\ -0.001753\pm 0.076580$ $0.129607\pm 0.026449$
Lite-CNN-Cons 1 $803315\pm 0.205572 0.0$$00106\pm 0.000029 0.0$$006285\pm 0.057660\ /\ -0.015335\pm 0.052686$ $0.120755\pm 0.029468$
ResNet18P-Cons 0 $439964\pm 0.247314 0.0$$74461\pm 0.065610 0.0$$14477\pm 0.047350\ /\ 0.000779\pm 0.038367 0.0$$048051\pm 0.056600$
ResNet18P-Cons 0.5 $920251\pm 0.196921 0.0$$06523\pm 0.003518 0.0$$06042\pm 0.012812\ /\ -0.002165\pm 0.015076 0.0$$067112\pm 0.031388$
ResNet18P-Cons 1 $931452\pm 0.163280 0.0$$03425\pm 0.001379 0.0$$02706\pm 0.028186\ /\ 0.009707\pm 0.019310 0.0$$095146\pm 0.057933$
Takeaways.
@@
Across all settings, we observe that pixel-space $\ell_{\infty}$ perturbations can severely degrade next-day direction prediction even at extremely small budgets. Robustness is strongly view-dependent: perturbing the indicator view is particularly damaging for ResNet18-P, often leading to near-saturated failure already at $\epsilon_{\mathrm{adv}}=0.25/255$. In the two-view regime, dual-branch late fusion consistently improves robustness under view-constrained attacks, supporting the intuition that an unperturbed branch can provide a partial evidence path. However, joint perturbations remain challenging and can still induce large performance drops, indicating that cross-view redundancy alone does not guarantee worst-case robustness. Within late-fusion models, increasing the consistency weight $\lambda$ yields secondary but measurable gains in several view-constrained scenarios, while the dominant robustness improvement stems from the architectural shift from early to late fusion.
Diagnostics. Branch diagnostics (Table 4, Appendix F) show that $\lambda$ mainly enforces logit-level alignment across views, while its impact on fused performance is secondary and backbone-dependent.
@@ 7 Discussion
Learnability hinges on the label-noise regime, with a non-monotonic data–noise trade-off.
A central observation from the clean baselines is that next-day direction prediction is near chance on the full dataset and becomes measurably learnable only on ex-post evaluation subsets defined by the minimum-movement threshold (min_move; Sec. 6.1), i.e., samples satisfying $|r_{t+1}|\geq\tau$. At small thresholds ($\tau\leq 0.004$), performance remains near chance across architectures, consistent with direction labels being dominated by micro-moves and label ambiguity. Increasing $\tau$ reduces near-zero ambiguity and can reveal predictive signal on the resulting filtered subset, but overly strict filtering ($\tau\approx 0.010$) reduces the effective sample size and increases seed-to-seed variance, leading to a non-monotonic regime in which additional “denoising” can harm reliability. We emphasize that min_move is a label-dependent subset definition rather than an inference-time filtering rule. This highlights a practical mechanism: in financial-image settings, reported generalization and robustness trends should be interpreted jointly with the evaluation regime induced by such subset definitions, rather than as properties of a fixed unconditional prediction task.
Fusion is noise-regime dependent: early fusion can induce negative transfer under high label noise.
The dual-view setting exhibits a clear regime shift. Under weak filtering, early fusion by channel stacking can underperform single-view models (and even become negative for Lite-CNN), suggesting that naively merging heterogeneous, noisy views encourages spurious correlations and negative transfer. Once labels stabilize ($\tau\geq 0.006$), dual-view learning becomes consistently beneficial and competitive, indicating that cross-view information is exploitable when the supervision signal is sufficiently reliable. In this stabilized regime, dual-branch late fusion provides a more reliable default than early fusion on clean evaluation, consistent with the large clean MCC gains observed at $\tau=0.006$. Overall, these results support a broader multimodal lesson: the relative advantage of multimodal fusion depends not only on model capacity, but also on the effective label-noise regime and the alignment between views.
Late fusion provides a robust evidence path under view-constrained corruption, but joint attacks remain challenging.
Across backbones, late fusion consistently improves robustness under view-constrained attacks (Sec. 6.2). A natural interpretation is that, when only one view is perturbed, the unperturbed branch can still provide partial evidence, preventing worst-case collapse and yielding flatter robustness curves over the budget grid. This robustness trend complements the clean-baseline finding that late fusion is a stronger and more reliable default than early fusion once labels stabilize. However, joint perturbations remain challenging: when both channels are attacked simultaneously, redundancy across views is insufficient to guarantee worst-case robustness, and performance can degrade substantially even for late-fusion models. This delineates the protection provided by multi-view redundancy: it helps against single-view corruption but does not substitute for robust training or stronger defenses against coordinated multi-view attacks.
View asymmetry reveals which information source is brittle under pixel-space perturbations.
A striking result is the strong view dependence of adversarial vulnerability, especially for ResNet18-P under indicator-view attacks where MCC can saturate near the lower bound at extremely small budgets (e.g., $\epsilon_{\mathrm{adv}}=0.25/255$), suggesting near-consistent label-flip behavior rather than a gradual attenuation of predictive signal. In contrast, OHLCV-view attacks often yield a more gradual deterioration. While both views are derived from the same underlying time series, their rendered representations and inductive biases differ, and the view-aligned evaluation protocol makes this asymmetry explicit. More broadly, view-aligned robustness curves can be interpreted as a diagnostic of which information source (or rendering) is brittle, guiding both representation choices and defense priorities.
Consistency regularization: logit-level alignment is easy; improved robustness is conditional.
Branch diagnostics (Table 4, Appendix F) help explain why consistency has modest, sometimes non-monotonic effects: increasing $\lambda$ rapidly tightens cross-branch logit agreement and reduces mismatch (sym_kl), yet branch-wise MCC stays near zero even when fused-head MCC is much higher, indicating that most predictive signal is carried by the fused late-fusion pathway rather than by standalone branch classifiers. For Lite-CNN, stronger consistency can over-align branch logits without improving discriminability, consistent with its non-monotonic clean and robustness trends across $\lambda$. For ResNet18-P, larger $\lambda$ improves fused performance in our setting but does not avert the saturated failure mode under indicator-view attacks. Although $\lambda=0.5$ is clean-optimal at $\tau=0.006$, we use $\lambda=1$ in robustness summaries because it yields slightly better robustness (including the $2/255$ stress test) at a small cost in clean MCC. Overall, consistency regularization is best viewed as a secondary stabilizer whose optimal strength is backbone- and regime-dependent, and should be tuned jointly with the fusion design and the label-noise operating point.
Threat-model interpretation and limitations.
Our main robustness evaluation adopts a pixel-space $\ell_{\infty}$ threat model on rendered financial images. This should be interpreted as a stress test of sensitivity to small input perturbations in the chosen representation, rather than a claim that real-world market adversaries directly manipulate pixels. Nevertheless, the observed fragility at tiny budgets underscores that these models can rely on brittle, non-robust features in the image parameterization. Several limitations follow. First, robustness conclusions are specific to the rendering and view–channel mapping used here; alternative parameterizations (e.g., different indicator encodings or chart styles) may change the attack surface. Second, the min_move filter couples label stability with sample size, so robustness comparisons should be contextualized by the induced data regime. Third, naturally trained models remain vulnerable to coordinated joint attacks; addressing worst-case robustness likely requires adversarial training, certified defenses, or representation-level constraints. Finally, our PGD evaluation uses a deterministic setting without random restarts; stronger first-order baselines (e.g., multi-restart PGD) and ensemble attacks (e.g., AutoAttack) are natural extensions for future work to further probe worst-case robustness.
Implications and future directions.
The results suggest several actionable directions for robustness and multi-view learning under noisy financial supervision. First, robustness evaluations should report the label-noise operating point explicitly (e.g., via $\tau$) and treat learnability as part of the experimental protocol. Second, in our dual-view setting, late fusion is a reliable starting point—especially under single-view corruption—though it does not guarantee worst-case robustness under joint perturbations. Third, diagnostics that separate logit-level alignment from fused predictive power can prevent over-interpreting agreement metrics as evidence of improved discriminability. Future work should complement pixel-space attacks with pipeline-level stress tests that reflect realistic variability—including small window-boundary misalignment, missingness/imputation effects, and rendering/normalization differences—followed by re-rendering the views. Benchmarks should also account for abrupt changes in market conditions that can shift the data distribution over short periods. Finally, beyond representation-level attacks, adversarial or corrupted data/pipeline inputs may degrade reliability; deployments should therefore include lightweight safeguards such as data provenance checks, cross-view integrity checks (alignment, missingness, and abnormal divergence), and drift/anomaly monitoring.
@@ 8 Conclusion
We investigated same-source multi-view learning and adversarial robustness for next-day direction prediction on SGE gold spot data (2005–2025), using two deterministically constructed and window-aligned image views: an OHLCV-rendered price/volume chart (ohlcv) and a technical-indicator matrix view (indic). This controlled setup isolates representation and fusion effects from cross-source alignment confounds.
Across leakage-resistant time-block splits with embargo, we find that both learnability and robustness trends depend strongly on the effective label-ambiguity regime induced by near-zero moves. We therefore use an ex-post minimum-movement filter (min_move), defined using the realized next-day return magnitude $|r_{t+1}|$, to construct evaluation subsets that exclude near-zero moves and reduce direction-label ambiguity. This induces a non-monotonic data–noise trade-off: moderate $\tau$ can make the conditional direction task on the resulting subset more learnable, while overly strict filtering reduces effective sample size and increases variance. Importantly, min_move is used for offline benchmark/subset definition and robustness analysis rather than as an inference-time decision rule. In the filtered subsets (default $\tau=0.006$), fusion is strongly regime dependent: early fusion by channel stacking can suffer negative transfer under noisier operating points, whereas late fusion with dual encoders and a fusion head is a more reliable default and provides the dominant clean-performance gains; cross-view consistency regularization yields secondary, backbone- and regime-dependent improvements.
From a security/robustness perspective, naturally trained models are highly vulnerable to test-time pixel-space $\ell_{\infty}$ perturbations on rendered inputs, even at extremely small budgets. Using an explicit view–channel mapping, we evaluated two multi-view threat scenarios: view-constrained attacks that perturb exactly one view and joint attacks that perturb both views. Robustness is markedly view dependent; for some backbones, indicator-view perturbations can drive MCC toward its minimum possible values under tiny budgets. Late fusion consistently improves robustness under view-constrained attacks by preserving an unperturbed evidence path, but joint perturbations remain challenging and can still cause substantial worst-case degradation, indicating that cross-view redundancy alone does not guarantee worst-case robustness. Branch diagnostics suggest that consistency regularization primarily enforces logit-level alignment, while predictive signal is largely expressed through the feature-level fusion pathway rather than strong standalone branch classifiers.
Overall, our results highlight that (i) reported robustness trends in same-source financial imaging are highly sensitive to evaluation design—in particular, leakage-resistant time-block splits with embargo for overlapping windows and explicit reporting of operating points, including any ex-post subset definitions used to reduce near-zero-move label ambiguity—so these choices should be stated clearly to support reliable robustness assessment; (ii) the benefits of multi-view fusion are operating-point dependent, and can shift from negative transfer to consistent gains as label ambiguity is reduced; and (iii) robustness evaluation for multi-view inputs should distinguish view-constrained from joint attacks to properly characterize the attack surface.

### v2

@@ 6.2 Adversarial Robustness
Protocol and summary metrics
@@
Summary of findings
At the default operating point $\tau=0.006$, we observe (i) rapid degradation under tiny pixel-space budgets, (ii) pronounced view dependence that is backbone- and fusion-dependent, and (iii) a robustness advantage of late fusion under view-constrained attacks, while robustness under joint (two-view) attacks remains challenging. The sensitivity study at $\tau=0.008$ in Appendix˜I confirms the structural trends in (i) and (iii), but indicates that finer-grained hyperparameter conclusions (e.g., the best $\lambda$ within late fusion) can shift with the operating point. In contrast, the direction of view asymmetry is primarily model- and fusion-dependent: ResNet18-P is consistently most vulnerable to indic-only perturbations, whereas Lite-CNN-Cons in late fusion is consistently more vulnerable to ohlcv-only perturbations under view-constrained attacks (Tables˜2 and 11).
@@
Overall vulnerability under tiny $\ell_{\infty}$ budgets
@@
View sensitivity and view-aligned evaluation
Using the view–channel mapping in Section˜3.3, we compare (i) single-view models attacked on their sole channel and (ii) two-view models attacked on the corresponding view channel (Section˜4.2). The results reveal strong view dependence, but its direction is backbone- and fusion-dependent. For ResNet18-P, perturbing the indicator view is consistently the most destructive at both $\tau=0.006$ and $\tau=0.008$, and very small budgets can already drive MCC close to the lower bound, suggesting near-saturated failure rather than a gradual attenuation of predictive signal. For Lite-CNN variants, view asymmetry is milder and depends on the fusion pathway; notably, for Lite-CNN-Cons with late fusion, ohlcv-only perturbations are consistently the most destructive single-view attacks at both $\tau$ values (Tables˜2 and 11).
Late fusion improves robustness under view-constrained attacks
A consistent trend in the two-view setting is that dual-branch late fusion (*-cons) degrades more gracefully than early fusion (channel stacking) under view-constrained attacks. When only one view is perturbed, the unperturbed branch can still provide usable evidence, which mitigates worst-case collapse and yields a flatter robustness curve over $E$. This advantage is most pronounced in the Lite-CNN family under single-view attacks (and persists at $\tau=0.008$; Appendix˜I). For ResNet18-P, late fusion improves robustness against OHLCV-only perturbations but does not prevent saturated failures under indicator-only attacks.
Joint perturbations remain challenging
@@
Consistency regularization: alignment, not a universal robustness knob
Within late-fusion models, the consistency weight $\lambda$ primarily enforces cross-branch logit alignment, while its impact on fused-head robustness is secondary and can be non-monotonic across operating points. We therefore treat $\lambda$ as a tunable refinement on top of the late-fusion design rather than a universally optimal setting; detailed branch diagnostics supporting this interpretation are provided in Appendix˜H.
@@
Table 2: PGD robustness summary at $\tau=0.006$ (min_move). We report MCC at two representative budgets ($0.25/255$ and $1/255$) and the worst-case drop $\Delta_{\mathrm{worst}}$ over nonzero budgets $\{0.25,0.5,0.75,1\}/255$ (Eq. (8)). As an additional stress test, results at $\epsilon_{\mathrm{adv}}=2/255$ are reported in Appendix˜J.
@@
Model Input Attack (PGD) MCC@0 MCC@0.25/255 MCC@1/255 $\boldsymbol{\Delta_{\mathrm{worst}}}$ Lite-CNN ohlcv on ohlcv 0.109 0.054 -0.088 -0.197 Lite-CNN indic on indic 0.079 0.024 -0.186 -0.265 Lite-CNN both (early) on ohlcv (ch0) 0.089 0.030 -0.104 -0.193 Lite-CNN both (early) on indic (ch1) 0.089 0.033 -0.094 -0.183 Lite-CNN both (early) on both 0.089 -0.014 -0.317 -0.406 Lite-CNN-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.121 0.081 -0.029 -0.150 Lite-CNN-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.121 0.113 0.084 -0.037 Lite-CNN-Cons ($\lambda{=}1$) both (late) on both 0.121 0.071 -0.072 -0.192 ResNet18-P ohlcv on ohlcv 0.019 -0.443 -0.952 -0.971 ResNet18-P indic on indic 0.082 -0.999 -1.000 -1.082 ResNet18-P-Cons ($\lambda{=}1$) both (late) on ohlcv (ch0) 0.095 -0.211 -0.774 -0.870 ResNet18-P-Cons ($\lambda{=}1$) both (late) on indic (ch1) 0.095 -0.959 -1.000 -1.095 ResNet18-P-Cons ($\lambda{=}1$) both (late) on both 0.095 -0.983 -1.000 -1.095
Takeaways
@@
Across all settings, pixel-space $\ell_{\infty}$ perturbations can severely degrade next-day direction prediction even at extremely small budgets. Robustness is strongly view-dependent and architecture-specific (ResNet18-P is particularly fragile to indic-only attacks). In the two-view regime, dual-branch late fusion generally improves robustness under view-constrained attacks, supporting the intuition that an unperturbed branch can provide a partial evidence path. However, joint perturbations remain challenging and can still induce large performance drops. Finally, consistency regularization mainly controls cross-branch alignment; its effect on robustness is secondary and can depend on the operating point (Appendix˜I).
PGD stress test at $\epsilon_{\mathrm{adv}}=2/255$ (not in the main grid)
We extend the perturbation budget to $\epsilon_{\mathrm{adv}}=2/255$ for late-fusion two-view Lite-CNN-Cons models and find that they remain non-trivially robust under the view-constrained indic-only attack (ch1), while ohlcv-only and joint attacks are substantially more destructive; full results at both $\tau=0.006$ and $\tau=0.008$ are reported in Appendix˜J.
@@ 7 Discussion
Operating point, noise regime, and fusion
All findings should be interpreted relative to the operating point induced by min_move ($\tau$) in Section˜6.1. Increasing $\tau$ reduces direction-label ambiguity by filtering near-zero moves but also reduces effective sample size and can increase variance. In noisy regimes (weak filtering), early fusion by channel stacking can underperform the stronger single-view baseline, consistent with negative transfer; once labels stabilize (around $\tau\geq 0.006$ in our setup), using both views becomes more reliably beneficial and dual-branch late fusion is a robust default. Sensitivity at $\tau=0.008$ in Appendix˜I preserves the qualitative patterns, while showing that some within-family hyperparameter preferences can shift with the operating point.
Robustness under tiny budgets: redundancy helps only under view-constrained attacks
Across backbones and operating points, naturally trained models can degrade sharply under small pixel-space $\ell_{\infty}$ budgets (Sections˜6.2, 2 and 11). Late fusion often degrades more gracefully under view-constrained attacks, consistent with preserving an unperturbed evidence path when only one view is perturbed. This mitigation is limited: view-specific failure modes can still saturate quickly, and joint perturbations remain challenging for both early and late fusion, including in the $2/255$ stress test (Appendix˜J).
Stable view asymmetry as a diagnostic
Vulnerability is strongly view dependent. ResNet18-P is consistently most fragile to indicator-view perturbations, while Lite-CNN-Cons under late fusion is more sensitive to ohlcv-only perturbations under view-constrained attacks. This suggests that robustness depends on view rendering and backbone inductive bias, and the direction of view asymmetry provides a useful diagnostic for interpreting view-aligned robustness results and identifying view-specific failure modes.
Consistency regularization and limitations
Branch diagnostics in Appendix˜H show that increasing $\lambda$ improves cross-branch agreement, but robustness effects are conditional: the clean-optimal $\lambda$ is backbone-dependent at $\tau=0.006$ (Section˜6.1), and robustness rankings among $\lambda\in\{0,0.5,1\}$ can change with $\tau$ (Appendix˜I). Our study evaluates pixel-space attacks on rendered images (a representation-level stress test), uses a label-dependent subset definition for offline benchmarking (Appendix˜D), and benchmarks only convolutional backbones; broader model families and threat models remain valuable extensions.
@@ 8 Conclusion
We studied same-source multi-view learning and adversarial robustness for next-day direction prediction using two deterministic, window-aligned image views (ohlcv and indic), enabling explicit view–channel threat modeling. Under leakage-resistant time-block splits with embargo, learnability and robustness depend strongly on the label-ambiguity regime induced by near-zero moves. We therefore use an ex-post minimum-movement filter (min_move) to define conditional evaluation subsets, revealing a non-monotonic data–noise trade-off: moderate $\tau$ improves learnability on the remaining subset, while overly strict filtering reduces sample size and increases variance. Importantly, min_move is a label-dependent subset definition for offline benchmarking and is not an inference-time decision rule (Appendix˜D).
In stabilized regimes (default $\tau=0.006$), early fusion can suffer negative transfer under noisier settings, whereas dual-branch late fusion is a more reliable default and provides the dominant clean-performance gains (Section˜6.1). From a robustness perspective, naturally trained models can be highly vulnerable to pixel-space $\ell_{\infty}$ perturbations even at tiny budgets (Section˜6.2). Late fusion often degrades more gracefully under view-constrained attacks, but joint perturbations remain challenging, indicating that cross-view redundancy alone does not guarantee worst-case robustness (Appendix˜J). Sensitivity at $\tau=0.008$ preserves the qualitative conclusions while highlighting operating-point dependence in within-family hyperparameters (Appendix˜I). Overall, reliable robustness assessment in same-source financial imaging benefits from leakage-resistant temporal evaluation, explicit reporting of operating points and label-dependent subset definitions, and threat models that distinguish view-constrained from joint multi-view attacks.
Declaration of competing interests
The author declares no competing interests.